OpenStudy (anonymous):

solve |x+3| +7 (is less than or equal to) 10

OpenStudy (anonymous):

First solve it for the absolute value part: \[\implies |x+3| \le 10-7\]\[\implies |x+3| \le 3\] Then recall that: \[|a| \le b \implies \cases{ \begin{array}{ccc}-b \le a \le b & \text{if} & b > 0\\ a = 0 & \text{if} & b = 0\\ \text{No solutions} & \text{if} & b < 0\end{array}}\]

OpenStudy (anonymous):

Since 3 is greater than 0.. \[\implies -3 \le x + 3 \le 3\]\[\implies -6 \le x \le 0\]

OpenStudy (anonymous):

Describe the graph of the solution in a complete sentence